# Modality Inflation: Energy Characterization and Optimization Opportunities for MLLM Inference

### v1

@@ Abstract.
Sustainable AI, Multimodal LLMs, Energy efficiency
1. Introduction
@@
Large language models (LLMs) have undergone rapid advances in recent years, and their evolution has shifted toward multimodal large language models (MLLMs) (wadekar2024evolution). By integrating text with additional modalities such as images and videos, MLLMs extend the capabilities of LLMs and have become increasingly important for applications (yin2024survey) including visual question answering, image captioning, and scene understanding (schwenk2022okvqa; chen2022visualgpt; sarto2025image; fan2024mllm). Despite their growing adoption, MLLMs differ from text-only LLMs in architectural complexity, workload characteristics, and resource demands. As illustrated in Figure 1, MLLMs for image-text requests introduce a modality-specific encoder alongside the LLM backbone, which not only increases computational requirements but also produces hundreds to thousands of visual tokens that propagate into the prefill stage and reshape the inference pipeline (chen2024image). This additional stage substantially raises overall energy consumption, making energy efficiency a critical concern for large-scale deployment.
@@
Meeting the escalating energy demands of MLLMs poses significant challenges for real-world serving systems. Prior work on text-only LLMs has shown that inference on modern GPUs is highly energy-intensive, with costs scaling superlinearly with model size and workload throughput (sarto2025image). These challenges are magnified in multimodal models, where vision encoders expand input complexity and inflate visual token counts (cai2024vip). Operating such energy-hungry workloads at datacenter scale raises sustainability concerns and operational costs, highlighting the need for energy-efficient serving strategies in large-scale MLLM deployments. To design energy-efficient systems, it is first necessary to develop a clear understanding of the models, workloads, and the sources of additional energy consumption for MLLMs.
@@
Figure 1. The MLLM inference pipeline. Multimodal inputs introduce a vision encoder and visual tokens (v), which are concatenated with text tokens (t) and inflate downstream prefill computation compared to text-only LLMs.
@@
Recent research has begun to address these challenges along two main directions. On the system side, several frameworks modularize MLLM architectures to enable more efficient serving and scheduling (qiumodserve; singh2024efficiently; dong2025hydrainfer; liu2025elasticmm). On the model side, techniques such as quantization, image compression, input sequence reduction, and token pruning reduce model size or limit the number of generated tokens, thereby lowering processing costs (wang2024q; song2024less; shang2024llava). However, while these approaches improve computation and throughput, they provide little insight into how energy consumption in MLLMs differs from text-only models or how it is distributed across inference stages, leaving opportunities for stage-wise energy optimization. This motivates our central objective to characterize how energy is distributed in multimodal pipelines and how input complexity propagates to higher energy cost.
@@
2. Background and Motivation
2.1. MLLM Architecture and Inference Pipeline
@@
MLLMs employ modality-specific encoders to transform non-text inputs, such as images or videos, into representations compatible with the LLM backbone. These encoders extract visual features and convert them into visual tokens that are aligned with the language model embedding space (yin2024survey). The inference pipeline typically consists of three stages: modality encoding, prefill, and decoding. In the encoding stage, the vision encoder generates visual tokens; this stage generally exhibits computational and memory behavior that lies between prefill and decoding (dong2025hydrainfer). The generated visual tokens are then concatenated with text tokens and processed during the prefill stage, whose cost scales with total sequence length and increases as visual tokens grow. Finally, decoding generates output tokens autoregressively while attending to the accumulating context window (qiumodserve). Compared to text-only inference, the additional encoding stage and the inflated input sequence introduce extra computation and memory overhead, fundamentally influencing the energy and performance characteristics of MLLM inference.
2.2. Encoder Architecture
@@
In MLLMs, the vision encoder plays a central role in shaping inference cost by determining how visual inputs are represented and how many tokens are exposed to the language model. Unlike text-only inference, the output size of this encoder varies widely across architectures, making encoder design a key factor in sequence length, memory footprint, and computational load. Existing MLLMs adopt different strategies for converting visual inputs into tokens. A common approach is fixed-patch tokenization, used in ViT-based models such as CLIP (radford2021learning), where images are divided into uniform patches to produce a constant token count (radford2021learning; liu2024improved). This yields predictable costs but does not scale with higher resolutions. In contrast, any-resolution tiling divides images into multiple crops or tiles, allowing flexible input handling at the cost of token counts that grow with resolution and aspect ratio (li2024llava; guo2024llava; luo2024feast). To mitigate token growth, some models compress visual features within the encoder through downsampling or lightweight attention (zhu2025internvl3; bai2025qwen2), while query-based architectures (e.g., Q-Former) project dense features into a fixed set of queries to bound token counts (li2023blip; kim2024towards). Additional mechanisms such as resampling, condensation, and token pruning further regulate token budgets (alayrac2022flamingo; li2025tokenpacker; shang2024llava).
2.3. Energy Saving Controls for GPUs
@@
Modern GPUs expose several mechanisms to balance performance and power during inference. A widely adopted control is Dynamic Voltage and Frequency Scaling (DVFS), which adjusts core or memory frequencies to reduce power consumption with limited performance impact (liu2025greenllm; maliakel2025investigating). Recent work shows that modest frequency reductions can yield noticeable energy savings for LLM workloads (10946802). Power capping offers another mechanism for managing GPU consumption by enforcing an upper power limit during execution. While effective for reducing peak power draw, the performance impact depends on workload characteristics and the severity of the cap. (patel2024characterizing). In this work, we focus on DVFS as a practical knob and evaluate its impact on energy efficiency during MLLM inference.
2.4. Modality Inflation Cost
@@
Although energy demand is already a growing concern for LLM inference (kim2024towards; kakolyris2024slo), multimodal workloads introduce an additional source of overhead by expanding the effective input sequence through visual tokens. This expansion increases both computation and memory traffic during inference and alters how energy is distributed across the pipeline. This modality-driven growth, which we refer to as modality inflation, increases both compute demand and memory traffic within the inference pipeline.
@@
2.5. Workload Heterogeneity in MLLM
There is a substantial degree of heterogeneity in multimodal requests. To better understand this behavior, we begin by analyzing the ServeGen (xiang2025servegen) trace of production-like queries and showing the cumulative distribution function of the number of images per query. As shown in Figure 2(a), many queries attach only one or two images. At the same time, we also observe queries with several dozen images, with rare extreme cases reaching up to 49 images per request. These image-heavy queries are relatively rare, but they still matter for the system. A small set of such requests is enough to slow down responses and increase resource usage, even though the average number of images per query is quite modest.
(a) Distribution of images per query from the ServeGen trace (xiang2025servegen).
@@ 2.5. Workload Heterogeneity in MLLM
Figure 2. Analysis of workload heterogeneity in multimodal requests.
We then place these findings in context by looking at image resolution across four widely used benchmark datasets for multimodal tasks: VQAv2 (goyal2017making), VizWiz (gurari2018vizwiz), ShareGPT4V (chen2024sharegpt4v), and ChartQA (masry2022chartqa). Figure 2(b) shows the cumulative distribution functions of image resolution for these datasets. The images range from small, low-resolution inputs to very large images with millions of pixels. The four curves are not identical. Some datasets are skewed toward smaller images, while others contain a larger fraction of high-resolution content. As a result, two queries that include the same number of images can still require very different amounts of visual-token computation, depending on which dataset or application they come from.
@@
3. Characterizing Energy and Power in MLLM Inference
@@
3.1. Experimental Setup
Models. We select four representative MLLMs, including LLaVA-1.5 (liu2024improved), LLaVA-OneVision(li2024llava), Qwen2.5-VL(bai2025qwen2), and InternVL3 (zhu2025internvl3), which span a range of vision encoder designs and tokenization strategies, as explained in Table 1. Together, these models represent common architectural choices in modern MLLMs, including fixed patch tokenization, any resolution tiling, compressed ViT based encoding, and query based token reduction. All models use LLM backbones in the 7B–8B parameter range, and three of them share the same LLM backbone architecture. This controlled selection allows us to isolate the impact of vision encoding and visual tokenization on energy consumption and keep the backbone scale comparable.
@@ 3.1. Experimental Setup
Hardware and Software. All experiments are conducted on a single NVIDIA A100 80GB GPU. Model inference is executed using the PyTorch (paszke2019pytorch) deep learning framework with Hugging Face Transformers (wolf2020transformers). To capture fine-grained energy profiles, we monitor GPU power consumption via the NVIDIA Management Library (NVML) (nvidia_nvml), sampling at 5 ms intervals. To ensure reproducibility, we allow utilization states to settle between consecutive runs and report averages over multiple repetitions with confidence intervals.
Model Vision Encoder LLM Backbone Average Accuracy (%) (duan2024vlmevalkit)
InternVL3-8B (zhu2025internvl3) InternViT-300M-v2.5 Qwen2.5-7B 73.6
LLaVA-1.5-7B (liu2024improved) CLIP ViT-L/14 Vicuna-v1.5-7B 36.9
LLaVA-OneVision-Qwen2-7B (li2024llava) SigLIP ViT Qwen2-7B 60.2
Qwen2.5-VL-7B (bai2025qwen2) QwenViT Qwen2.5-7B 70.9
Table 1. Architecture overview of the MLLMs evaluated in this study, including vision encoders, LLM backbones, and representative average accuracy reported in prior work.
3.2. Energy Overhead of MLLMs
@@
Notably, energy and latency do not increase proportionally in this controlled setting. For Qwen2.5-VL, a 94% energy increase corresponds to a 179% latency increase, suggesting that a low parallelism stage can dominate end-to-end time. Overall, these results indicate that, unlike text-only LLM inference where energy behavior is relatively consistent across similarly sized models (pronk2025benchmarking), MLLMs can exhibit large architecture dependent differences even under controlled input conditions, motivating model specific serving policies rather than one-size-fits-all configurations.
@@
Figure 3. Iso-token comparison of (a) energy per request and (b) latency per request between multimodal models and their text-only baselines. The total input token count is matched by equating text-only tokens with the combined text and visual tokens in the multimodal setting, and output generation is fixed to one token. Numeric labels inside the LLM bars indicate the number of tokens used for iso-token matching.
3.3. Stage-wise Characterization
@@
The latency breakdown also highlights different optimization opportunities: for LLaVA-1.5, the encoding stage is short (about 12 ms), suggesting potential latency slack for energy-saving techniques, whereas for LLaVA-OneVision the end-to-end latency is dominated by prefill, making this stage the primary target under tight response-time constraints. Across all models, decoding remains comparatively stable, with energy largely governed by output length rather than encoder design or input length (fernandez2025energy). These results show that energy bottlenecks vary across stages depending on the architecture, motivating stage-specific configurations and stage-wise DVFS.
@@
Figure 4. Stage-wise breakdown of inference into encoding, prefill, and decoding with a fixed output length, showing (a) energy per request and (b) latency per request for each stage. Numeric labels inside the prefill bars indicate the number of visual tokens produced by each model in this setting.
3.4. Power Profiling
@@
Figure 5. Normalized GPU power traces for text-only LLM and multimodal LLM inference using batch size 32 and output length 32 for GPU saturation.
To answer RQ3, we analyze instantaneous GPU power draw during inference by comparing normalized power traces of multimodal requests to iso-token text-only baselines. For these experiments, we use a batch size of 32 to represent a throughput optimized setting that drives the A100 toward saturation. As shown in Figure 5, the text-only baseline transitions rapidly from idle (about 80 W) to near the GPU power limit (about 400 W), indicating sustained high GPU activity in this optimized setting. In contrast, multimodal inference exhibits extended mid power phases, typically between 100 W and 250 W, which coincide with the vision encoding portion of the pipeline. Qwen2.5-VL shows the most pronounced effect, with a step-like power pattern around 200 W for a significant duration before the LLM backbone engages, while LLaVA-OneVision exhibits higher frequency fluctuations consistent with bursty tile processing. Importantly, while disaggregated serving can increase encoder batch size and improve utilization (qiumodserve; dong2025hydrainfer), such batching is constrained by stochastic request arrivals and end-to-end latency targets; therefore, mid power phases can still arise under realistic operating points. Overall, these trace signatures indicate that multimodal execution spends substantial time in a lower power regime than the saturated text-only baseline, suggesting that generic race-to-idle frequency policies may keep the GPU at high clocks during phases that do not benefit proportionally from higher frequency.
@@
Figure 6. Impact of image count on energy and latency in MLLM inference.
3.5. Input Complexity
To address RQ4, we characterize how energy and latency scale with two primary dimensions of multimodal input complexity, image resolution and image count, while keeping the text prompt fixed. This analysis is important for serving because real workloads vary widely in both resolution and the number of images per request, as discussed in Section 2.5, and worst-case inputs can dominate both energy consumption and tail latency. As shown in Figure 7(a) and 7(b), both energy and latency exhibit clearly different scaling behaviors across architectures. LLaVA-1.5 remains nearly flat as resolution increases, consistent with its fixed-resolution CLIP encoder that resizes inputs before tokenization. InternVL3 and LLaVA-OneVision exhibit non-uniform scaling over the evaluated resolution range, with noticeable increases at specific resolution settings. These trends are explained by how each model expands images into visual tokens. Figure 7(c) shows that LLaVA-1.5 produces an almost constant number of visual tokens across resolutions, whereas InternVL3 and LLaVA-OneVision increase token counts discretely over the tested resolutions, and Qwen2.5-VL exhibits rapid token growth at higher resolutions. In particular, Qwen2.5-VL shows a sharp increase in both energy and latency beyond $1024\times 1024$, and the accompanying token expansion suggests that high-resolution inputs can substantially increase prefill workload when token growth is not effectively controlled. Beyond resolution, Figure 6 shows that increasing the number of images increases both energy and latency over the evaluated range, but with markedly different slopes across models. These slopes capture the marginal cost per additional image and directly constrain multi-image serving under fixed response-time targets.

### v2

@@ Abstract
I Introduction
@@
Large language models (LLMs) have undergone rapid advances in recent years, and their evolution has shifted toward multimodal large language models (MLLMs) [38]. By integrating text with additional modalities such as images and videos, MLLMs extend the capabilities of LLMs and have become increasingly important for applications [43] including visual question answering, image captioning, and scene understanding [33, 4, 32, 8]. Despite their growing adoption, MLLMs differ from text-only LLMs in architectural complexity, workload characteristics, and resource demands. As illustrated in Figure 1, MLLMs for image-text requests introduce a modality-specific encoder alongside the LLM backbone, which not only increases computational requirements but also produces hundreds to thousands of visual tokens that propagate into the prefill stage and reshape the inference pipeline [5]. This additional stage substantially raises overall energy consumption, making energy efficiency a critical concern for large-scale deployment.
@@
Meeting the escalating energy demands of MLLMs poses significant challenges for real-world serving systems. Prior work on text-only LLMs has shown that inference on modern GPUs is highly energy-intensive, with costs scaling superlinearly with model size and workload throughput [25]. These challenges are magnified in multimodal models, where vision encoders expand input complexity and inflate visual token counts [3]. Operating such energy-hungry workloads at datacenter scale raises sustainability concerns and operational costs, highlighting the need for energy-efficient serving strategies in large-scale MLLM deployments. To design energy-efficient systems, it is first necessary to develop a clear understanding of the models, workloads, and the sources of additional energy consumption for MLLMs.
@@
Figure 1: The MLLM inference pipeline. Multimodal inputs introduce a vision encoder and visual tokens (v), which are concatenated with text tokens (t) and inflate downstream prefill computation compared to text-only LLMs.
@@
Recent research has begun to address these challenges along two main directions. On the system side, several frameworks modularize MLLM architectures to enable more efficient serving and scheduling [30, 35, 7, 21]. On the model side, techniques such as quantization, image compression, input sequence reduction, and token pruning reduce model size or limit the number of generated tokens, thereby lowering processing costs [39, 36, 34]. However, while these approaches improve computation and throughput, they provide little insight into how energy consumption in MLLMs differs from text-only models or how it is distributed across inference stages, leaving opportunities for stage-wise energy optimization. This motivates our central objective to characterize how energy is distributed in multimodal pipelines and how input complexity propagates to higher energy cost.
@@
II Background and Motivation
II-A MLLM Architecture and Inference Pipeline
@@
MLLMs employ modality-specific encoders to transform non-text inputs, such as images or videos, into representations compatible with the LLM backbone. These encoders extract visual features and convert them into visual tokens that are aligned with the language model embedding space [43]. The inference pipeline typically consists of three stages: modality encoding, prefill, and decoding. In the encoding stage, the vision encoder generates visual tokens; this stage generally exhibits computational and memory behavior that lies between prefill and decoding [7]. The generated visual tokens are then concatenated with text tokens and processed during the prefill stage, whose cost scales with total sequence length and increases as visual tokens grow. Finally, decoding generates output tokens autoregressively while attending to the accumulating context window [30]. Compared to text-only inference, the additional encoding stage and the inflated input sequence introduce extra computation and memory overhead, fundamentally influencing the energy and performance characteristics of MLLM inference.
II-B Encoder Architecture
@@
In MLLMs, the vision encoder plays a central role in shaping inference cost by determining how visual inputs are represented and how many tokens are exposed to the language model. Unlike text-only inference, the output size of this encoder varies widely across architectures, making encoder design a key factor in sequence length, memory footprint, and computational load. Existing MLLMs adopt different strategies for converting visual inputs into tokens. A common approach is fixed-patch tokenization, used in ViT-based models such as CLIP, where images are divided into uniform patches to produce a constant token count [31, 19]. This yields predictable costs but does not scale with higher resolutions. In contrast, any-resolution tiling divides images into multiple crops or tiles, allowing flexible input handling at the cost of token counts that grow with resolution and aspect ratio [16, 12, 22]. To mitigate token growth, some models compress visual features within the encoder through downsampling or lightweight attention [44, 2], while query-based architectures (e.g., Q-Former) project dense features into a fixed set of queries to bound token counts [17, 15]. Additional mechanisms such as resampling, condensation, and token pruning further regulate token budgets [1, 18, 34].
II-C Energy Saving Controls for GPUs
@@
Modern GPUs expose several mechanisms to balance performance and power during inference. A widely adopted control is Dynamic Voltage and Frequency Scaling (DVFS), which adjusts core or memory frequencies to reduce power consumption with limited performance impact [20, 23]. Recent work shows that modest frequency reductions can yield noticeable energy savings for LLM workloads [37]. Power capping offers another mechanism for managing GPU consumption by enforcing an upper power limit during execution. While effective for reducing peak power draw, the performance impact depends on workload characteristics and the severity of the cap [28]. In this work, we focus on DVFS as a practical knob and evaluate its impact on energy efficiency during MLLM inference.
II-D Modality Inflation Cost
@@
Although energy demand is already a growing concern for LLM inference [15, 14], multimodal workloads introduce an additional source of overhead by expanding the effective input sequence through visual tokens. This expansion increases both computation and memory traffic during inference and alters how energy is distributed across the pipeline. This modality-driven growth, which we refer to as modality inflation, increases both compute demand and memory traffic within the inference pipeline.
@@
II-E Workload Heterogeneity in MLLM
There is a substantial degree of heterogeneity in multimodal requests. To better understand this behavior, we begin by analyzing the ServeGen [42] trace of production-like queries and showing the cumulative distribution function of the number of images per query. As shown in Figure 2(a), many queries attach only one or two images. At the same time, we also observe queries with several dozen images, with rare extreme cases reaching up to 49 images per request. These image-heavy queries are relatively rare, but they still matter for the system. A small set of such requests is enough to slow down responses and increase resource usage, even though the average number of images per query is quite modest.
(a) Distribution of images per query from the ServeGen trace [42].
@@ II-E Workload Heterogeneity in MLLM
Figure 2: Analysis of workload heterogeneity in multimodal requests.
We then place these findings in context by looking at image resolution across four widely used benchmark datasets for multimodal tasks: VQAv2 [10], VizWiz [13], ShareGPT4V [6], and ChartQA [24]. Figure 2(b) shows the cumulative distribution functions of image resolution for these datasets. The images range from small, low-resolution inputs to very large images with millions of pixels. The four curves are not identical. Some datasets are skewed toward smaller images, while others contain a larger fraction of high-resolution content. As a result, two queries that include the same number of images can still require very different amounts of visual-token computation, depending on which dataset or application they come from.
@@
III Characterizing Energy and Power in MLLM Inference
Model Vision Encoder LLM Backbone Avg. Acc. (%)
InternVL3-8B [44] InternViT-300M-v2.5 Qwen2.5-7B 73.6
LLaVA-1.5-7B [19] CLIP ViT-L/14 Vicuna-v1.5-7B 36.9
LLaVA-OneVision-Qwen2-7B [16] SigLIP ViT Qwen2-7B 60.2
Qwen2.5-VL-7B [2] QwenViT Qwen2.5-7B 70.9
TABLE I: Architecture overview of the MLLMs evaluated in this study.
@@ III Characterizing Energy and Power in MLLM Inference
III-A Experimental Setup
Models. We select four representative MLLMs, including LLaVA-1.5 [19], LLaVA-OneVision[16], Qwen2.5-VL[2], and InternVL3 [44], which span a range of vision encoder designs and tokenization strategies, as explained in Table I. Together, these models represent common architectural choices in modern MLLMs, including fixed-patch tokenization, any-resolution tiling, compressed ViT-based encoding, and query-based token reduction. All models use LLM backbones in the 7B–8B parameter range, and three of them share the same LLM backbone architecture. This controlled selection allows us to isolate the impact of vision encoding and visual tokenization on energy consumption and keep the backbone scale comparable.
@@ III-A Experimental Setup
Hardware and Software. All experiments are conducted on a single NVIDIA A100 80GB GPU. Model inference is executed using the PyTorch [27] deep learning framework with Hugging Face Transformers [41]. To capture fine-grained energy profiles, we monitor GPU power consumption via the NVIDIA Management Library (NVML) [26], sampling at 5 ms intervals. To ensure reproducibility, we allow utilization states to settle between consecutive runs and report averages over multiple repetitions with confidence intervals.
III-B Energy Overhead of MLLMs
@@
Notably, energy and latency do not increase proportionally in this controlled setting. For Qwen2.5-VL, a 94% energy increase corresponds to a 179% latency increase, suggesting that a low parallelism stage can dominate end-to-end time. Overall, these results indicate that, unlike text-only LLM inference where energy behavior is relatively consistent across similarly sized models [29], MLLMs can exhibit large architecture dependent differences even under controlled input conditions, motivating model specific serving policies rather than one-size-fits-all configurations.
@@
Figure 3: Iso-token comparison of (a) energy per request and (b) latency per request between multimodal models and their text-only baselines. The total input token count is matched by equating text-only tokens with the combined text and visual tokens in the multimodal setting, and output generation is fixed to one token. Numeric labels inside the LLM bars indicate the number of tokens used for iso-token matching.
III-C Stage-wise Characterization
@@
The latency breakdown also highlights different optimization opportunities: for LLaVA-1.5, the encoding stage is short (about 12 ms), suggesting potential latency slack for energy-saving techniques, whereas for LLaVA-OneVision the end-to-end latency is dominated by prefill, making this stage the primary target under tight response-time constraints. Across all models, decoding remains comparatively stable, with energy largely governed by output length rather than encoder design or input length [9]. These results show that energy bottlenecks vary across stages depending on the architecture, motivating stage-specific configurations and stage-wise DVFS.
@@
Figure 4: Stage-wise breakdown of inference into encoding, prefill, and decoding with a fixed output length, showing (a) energy per request and (b) latency per request for each stage. Numeric labels inside the prefill bars indicate the number of visual tokens produced by each model in this setting.
III-D Power Profiling
@@
Figure 5: Normalized GPU power traces for text-only LLM and multimodal LLM inference using batch size 32 and output length 32 to drive the GPU toward saturation.
To answer RQ3, we analyze instantaneous GPU power draw during inference by comparing normalized power traces of multimodal requests to iso-token text-only baselines. For these experiments, we use a batch size of 32 to represent a throughput optimized setting that drives the A100 toward saturation. As shown in Figure 5, the text-only baseline transitions rapidly from idle (about 80 W) to near the GPU power limit (about 400 W), indicating sustained high GPU activity in this optimized setting. In contrast, multimodal inference exhibits extended mid power phases, typically between 100 W and 250 W, which coincide with the vision encoding portion of the pipeline. Qwen2.5-VL shows the most pronounced effect, with a step-like power pattern around 200 W for a significant duration before the LLM backbone engages, while LLaVA-OneVision exhibits higher frequency fluctuations consistent with bursty tile processing. Importantly, while disaggregated serving can increase encoder batch size and improve utilization [30, 7], such batching is constrained by stochastic request arrivals and end-to-end latency targets; therefore, mid power phases can still arise under realistic operating points. Overall, these trace signatures indicate that multimodal execution spends substantial time in a lower power regime than the saturated text-only baseline, suggesting that generic race-to-idle frequency policies may keep the GPU at high clocks during phases that do not benefit proportionally from higher frequency.
@@
Figure 6: Impact of image count on energy and latency in MLLM inference.
III-E Input Complexity
To address RQ4, we characterize how energy and latency scale with two primary dimensions of multimodal input complexity, image resolution and image count, while keeping the text prompt fixed. This analysis is important for serving because real workloads vary widely in both resolution and the number of images per request, as discussed in Section II-E, and worst-case inputs can dominate both energy consumption and tail latency. As shown in Figure 7(a) and 7(b), both energy and latency exhibit clearly different scaling behaviors across architectures. LLaVA-1.5 remains nearly flat as resolution increases, consistent with its fixed-resolution CLIP encoder that resizes inputs before tokenization. InternVL3 and LLaVA-OneVision exhibit non-uniform scaling over the evaluated resolution range, with noticeable increases at specific resolution settings. These trends are explained by how each model expands images into visual tokens. Figure 7(c) shows that LLaVA-1.5 produces an almost constant number of visual tokens across resolutions, whereas InternVL3 and LLaVA-OneVision increase token counts discretely over the tested resolutions, and Qwen2.5-VL exhibits rapid token growth at higher resolutions. In particular, Qwen2.5-VL shows a sharp increase in both energy and latency beyond $1024\times 1024$, and the accompanying token expansion suggests that high-resolution inputs can substantially increase prefill workload when token growth is not effectively controlled. Beyond resolution, Figure 6 shows that increasing the number of images increases both energy and latency over the evaluated range, but with markedly different slopes across models. These slopes capture the marginal cost per additional image and directly constrain multi-image serving under fixed response-time targets.